Laboratory Journal from 2003-04-22 to …

Jane Eager August 11, 2003

Index

Expression of Hagridin in E. coli, 3

Transformation with HagIV plasmid, 3

Lyophilization of A, 7 Lyophilization of B, 7

Preparation of A, 5 Preparation of B, 7 Preparation of purified something, 5 Preparation of RNase A S-peptide, 1–2,

5 subtilisin digest, 1–2, 5 Index iv

Contents

Tuesday, April 22, 2003 1

1 Preparation of RNase A S-peptide . . . . . . . . . . . . . . . . . . . . . . 1

1.1 subtilisin digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Expression of Hagridin in E. coli . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Transformation with HagIV plasmid . . . . . . . . . . . . . . . . . 3

Wednesday, April 23, 2003 5

1 Preparation of RNase A S-peptide . . . . . . . . . . . . . . . . . . . . . . 5

1.1 subtilisin digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2 Preparation of purified something . . . . . . . . . . . . . . . . . . . . . . . 5 3 Preparation of A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Thursday, April 24, 2003 7 1 Preparation of B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2 Lyophilization of A and B . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Contents vi

Tuesday, April 22, 2003

1 Preparation of RNase A S-peptide

1.1 Digestion with Subtilisin

A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges.

A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges.

Tuesday, April 22, 2003

A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges.

A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges.

Figure 1: Some very interesting SDS gel

2 Expression of Hagridin in E. coli

2.1 Transformation of hogwarts-auxotrophic E. coli with the HagIV plasmid

Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Tuesday, April 22, 2003 4

Wednesday, April 23, 2003

1 Preparation of RNase A S-peptide 1.1 Digestion using the new batch of Subtilisin 2 Preparation of purified something 3 Preparation of A

A was prepared according to the protocol in Muller et al. and was temporarily stored in the freezer.

Wednesday, April 23, 2003 6

Thursday, April 24, 2003

1 Preparation of B

B was prepared according to the protocol by Smith and Fox.

2 Lyophilization of A and B